QUESTION 1 How much work does the electric fuld do in meng a-758 vc charge from gound (vsovn Do a Point whose poeertal is 6TvNgher? QUESTION 2 How much work does se electric field do in moving a proton from a point with a potentia of +1224v to a point where it is -50 av? Express your answer in joules Note. The charge on a proron is a 6alio QUESTION 3 anong is the electic seld between two parallel plates 0.00am apart if the potential dretence between them is 26gv? Thr phrate hou trong is essentially asking for the magnitude Aino Pay attention othe ainernate units this chapter use field QUESTION 4 The wieetre seid between two paralel plates connected to a 45v battery is 1 rvim How far apart of ae the plates? QUESTION 5
Answer #1

here,

1)

charge , q = - 7.58 * 10^-6 C

potential difference , dV = 67 - 0 V = 67 V

the work done by electric feild , W = q * V

W = - 7.58 * 10^-6 * 67 J

W = - 5.08 * 10^-4 J

the work done by electric feild is - 5.08 * 10^-4 J